OpenStudy (anonymous):

Last one :) 4x+3y=7 2x-7y=-6 thanks for those who helped. :) Specially Josephniel14 :D

OpenStudy (anonymous):

Hahah. No Prob. XD

OpenStudy (anonymous):

Answer Please? :( teehee:D

OpenStudy (anonymous):

lol, 4x + 3y = 7 -2(2x - 7y) = -6*- 2 4x +3y=7 -4x+14y=12 17y= 19 => y=19/17 4x + 3(19/17)= 7 4x + 57/17= 7 4x= 7 - 57/17= 119/17 - 57/17= 62/17 x= 62/17*4= 18/17

OpenStudy (anonymous):

oops mistake,,, x= 31/34

OpenStudy (anonymous):

Same procedure as I just explained in the other one you posted, multiply the second equation by 2 so that you have 4x in both, then substitute for one of the 4x's with the value of the other to compute x, etc etc

OpenStudy (anonymous):

right. thankyou ;)

OpenStudy (anonymous):

Sorry. I was supposed to answer your question until my computer got loco. Hahah. Sorry.